# The DREAMS Project: Disentangling the Impact of Halo-to-Halo Variance and Baryonic Feedback on Milky Way Dark Matter Density Profiles

### v1

@@ Abstract
Astrophysical searches for dark matter in the Milky Way require a reliable model for its density distribution, which in turn depends on the influence of baryonic feedback on the Galaxy. In this work, we utilize a new suite of Milky Way-mass halos from the DREAMS Project, simulated with Cold Dark Matter (CDM),to quantify the influence of baryon feedback and intrinsic halo-to-halo variance on dark matter density profiles. Our suite of 1024 halos varies over supernova and black hole feedback parameters from the IllustrisTNG model, as well as variations in two cosmological parameters. We find that Milky Way-mass dark matter density profiles in the IllustrisTNG model are largely insensitive to astrophysics and cosmology variations, with the dominant source of scatter instead arising from halo-to-halo variance. However, most of the (comparatively minor) feedback-driven variations come from the changes to supernova prescriptions. By comparing to dark matter-only simulations, we find that the strongest supernova wind energies are so effective at preventing galaxy formation that the halos are nearly entirely collisionless dark matter. Finally, regardless of physics variation, all the DREAMS halos are roughly consistent with a halo contracting adiabatically from the presence of baryons, unlike models that have bursty stellar feedback. This work represents a step toward assessing the robustness of Milky Way dark matter profiles, with direct implications for dark matter searches where systematic uncertainty in the density profile remains a major challenge.
@@
1 Introduction
Baryons make up only a small fraction of the total mass budget of the Universe (Planck Collaboration et al. 2016), yet they play a central role in driving galaxy evolution. Unlike the dark matter component, which is assumed to be cold and collisionless in the standard Cold Dark Matter (CDM) model, the baryonic component experiences interactive forces and energy dissipation. These interactions ultimately facilitate the condensation of gas at the centers of halos and the formation of the galaxies we observe (Fall & Efstathiou, 1980; Blumenthal et al., 1986). Once a galaxy is formed, baryons can impart feedback in their environments via various processes including supernovae, stellar winds, and active galactic nuclei (AGN) activity (e.g., Dekel & Silk, 1986; Larson, 1974; Somerville & Davé, 2015). Through interactions, dissipation, and feedback, baryons drive time-varying changes in the gravitational potential that shape galaxy evolution.
The role that baryons play can be clearly seen in the distribution of matter within a halo. Early dark matter-only (DMO) simulations showed that halos ubiquitously form a two-component power-law density profile known as the Navarro, Frenk, & White (1997; NFW) profile such that
@@ 1 Introduction
Simulations which self-consistently model the baryonic component provide one possible resolution to these tensions (Sales et al., 2022). A cored inner density can be achieved using sufficiently “bursty” supernova feedback (Chan et al., 2015; Lazar et al., 2020; Mostow et al., 2024). Repetitively cycling between slow/smooth inflows followed by fast/impulsive outflows creates central potential fluctuations that disrupt the orbits of dark matter particles and leads to decreased central densities (Governato et al., 2010; Pontzen & Governato, 2012; Oñorbe et al., 2015; Jahn et al., 2023, although this effect is mass-dependent with it being most pronounced in dwarf galaxies). However, even at fixed physics prescriptions, the diversity of halo assembly histories introduces substantial halo-to-halo variance in the resulting dark matter profiles (e.g., Abadi et al., 2010; Duffy et al., 2010; Di Cintio et al., 2014; Chan et al., 2015; Fitts et al., 2017; Lazar et al., 2020; Farahi et al., 2022). Disentangling this intrinsic variance from systematic shifts due to baryonic physics is therefore critical for interpreting both simulations and observations. Moreover, the level to which feedback is generically bursty and its evolutionary consequences across different galaxy populations remains uncertain (see, e.g., work by Hartley & Ricotti, 2016; Bhagwat et al., 2024; Garcia et al., 2023, 2024a, 2024b, 2025a, 2025b, 2025c).
@@
Current cosmological simulations of galaxy evolution follow several prescriptions for modeling the physics of baryons (see, e.g., Vogelsberger et al., 2020; Crain & van de Voort, 2023; Feldmann & Bieri, 2025, for reviews). These prescriptions often rely on “subgrid” models for the numerical implementations of unresolved physics. For example, the interstellar medium (ISM) is often treated with an effective equation of state that sets the behavior of dense ($n_{\rm H}\gtrsim 0.1~{\rm cm}^{-3}$) star forming gas (Springel & Hernquist, 2003; Schaye & Dalla Vecchia, 2008). Subgrid models require assumptions and oftentimes contain tunable parameters within the model. As a result, different simulation models make different predictions for the role of baryons and their associated feedback—even with relatively similar implementations (e.g., Chua et al., 2019, 2022; Garcia et al., 2024a, 2025a, 2025b). Some work has been done to quantify the variation in model predictions based on the input physics (e.g., Duffy et al., 2010; Torrey et al., 2014; Pillepich et al., 2018a; Chua et al., 2019, 2022; Font et al., 2020; Anbajagane et al., 2022; Kugel et al., 2023), but the number of model variations is usually few (on the order of $\sim 10$s of model variations). Recent simulation efforts such as Cosmology and Astrophysics with MachinE Learning Simulations (CAMELS; Villaescusa-Navarro et al., 2021; Ni et al., 2023) and DaRk MattEr and Astrophysics with Machine learning and Simulations Project (DREAMS; Rose et al., 2025a) have begun to more systematically explore the parameter space of well-tested simulation models via thousands of model variations. These efforts allow for quantification of parameter uncertainties while also providing large enough samples to begin quantifying the intrinsic halo-to-halo variance.
@@
One model that departs sharply from the standard approach is Feedback In Realistic Environments (FIRE; Hopkins et al., 2014, 2018), which explicitly resolves the multiphase interstellar medium. While FIRE has its own subgrid assumptions and parameters, the treatment of the star forming ISM is a marked improvement compared to the standard approach of an effective equation of state. One feature of the FIRE model is the production of strong, time-variable stellar feedback-driven winds (i.e., bursty feedback; Muratov et al., 2015), whereas models that adopt effective equation of state ISM treatment yield smoother, less variable stellar feedback. In the absence of this bursty feedback, the evolution of a dark matter halo is governed by the response of the dark matter to the baryons cooling to the center. This adiabatic contraction can be modeled via the conservation of action integrals (e.g., Eggen et al., 1962; Blumenthal et al., 1986; Gnedin et al., 2004; Cautun et al., 2020): as the potential deepens due to baryon mass accumulation, the radius of the dark matter orbits decreases causing the total density in the center to increase. Feedback can partially offset or even reverse this contraction, but the degree depends sensitively on feedback (Brooks & Zolotov, 2014; Lovell et al., 2018; Hussein et al., 2025). Indeed, Hussein et al. (2025) show that FIRE-2 galaxies deviate much more from adiabatic contraction than simulations with equation of state ISM prescriptions, though it remains unclear whether this owes specifically to burstiness or just the overall feedback strength.
@@
The rest of this paper is organized as follows. In Section 2, we introduce the DREAMS simulations, their parameter variations, and halo selection. We then outline our method for reconstructing dark matter density fields, describe the neural network emulators used in this work, and discuss a scheme to account for how well our halos match observed galaxy scaling relations. In Section 3, we discuss the dark matter density profiles and fit them with an analytic function. In Section 4, we compare the hydrodynamic simulations to DMO simulations, quantify the contraction the halos must have undergone due to the presence of baryons, and contrast our results to those from the bursty feedback FIRE model. In Section 5, we state our conclusions. Finally, we present an example validation of our neural network emulation in Appendix A, show the lack of sensitivity of our results to the adopted halo weighting scheme in Appendix B, investigate the AGN feedback variations in more detail in Appendix C, and discuss our adiabatic contraction fitting methods in Appendix D.
2 Methods
This section is organized as follows: we begin by describing the new DREAMS CDM suite (§2.1), we then describe how we reconstruct and fit the dark matter density profiles (§2.2), we introduce our neural network emulation scheme (§2.3), and, finally, we summarize the Rose et al. (2025b) halo weighting scheme we apply to our sample (§2.4).
2.1 DREAMS Simulations
@@
Table 1: Astrophysical and Cosmological Parameter Variations. The astrophysical ($\bar{e}_{w}$, $\kappa_{w}$, $\epsilon_{f,\,\mathrm{high}}$) and cosmological ($\Omega_{\rm M}$, $\sigma_{8}$) parameters varied in the CDM DREAMS Milky Way-mass zoom-in simulations. We list the fiducial TNG value for each parameter, the variations used in these simulations (corresponding to a factor of 44 for $\bar{e}_{w}$ and $\epsilon_{f,\,\mathrm{high}}$, a factor of 22 for $\kappa_{w}$, and the two-standard-deviation uncertainty on Planck Collaboration et al. 2014 for $\Omega_{\rm M}$ and $\sigma_{8}$). We sample the unique set of parameters for each simulation according to a Sobol’ (1967) sequence. Finally, we note that $\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,\mathrm{high}}$ are sampled in logarithmic space, whereas $\Omega_{\rm M}$ and $\sigma_{8}$ are sampled in linear space. Detailed justifications for these parameter ranges can be found in Rose et al. (2025a).
This work makes use of data products from the DREAMS suites of simulations (Rose et al., 2025a).111dreams-project.org In particular, we use the zoom-in simulations from the CDM suite (Rose et al., 2025b). Each target halo falls within the mass range $5\times 10^{11}~{\rm M}_{\odot}<M_{\rm Halo}<2.0\times 10^{12}~{\rm M}_{\odot}$, roughly corresponding to the mass of the Milky Way halo (Callingham et al., 2019; Wang et al., 2020), where $M_{\rm Halo}$ is the sum of all mass within $R_{200}$ (Springel et al. 2005). There are a total of 2048 simulations in the DREAMS Milky Way CDM suite: 1024 DMO simulations as well as a corresponding pair of 1024 simulations with baryon physics included. Each pair has the same initial conditions, such that the same halo is simulated twice.
@@
The hydrodynamic simulations are based on the IllustrisTNG (hereafter simply TNG) physics model (Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018; Pillepich et al., 2018b; Springel et al., 2018; Pillepich et al., 2019; Nelson et al., 2019a, b). TNG is built on the moving Voronoi mesh code arepo (Springel, 2010a). The TNG model implements a wide range of astrophysical processes including (but not limited to) feedback from stars, the growth of supermassive black holes, feedback from AGN, cosmological expansion of the universe, gravity, dark matter, and galactic winds; we refer the reader to Pillepich et al. (2018a) for a complete description of the TNG model.
@@
The DREAMS halos are selected using an iterative zoom-in process (described in detail in Rose et al., 2025a, b). Briefly, the halos are selected initially from a low-resolution $(100~{\rm Mpc}/h)^{3}$ DMO simulation and are then resimulated with DMO zooms. The target halos of the intermediate zooms are selected to have masses of $5\times 10^{11}~{\rm M}_{\odot}<M_{\rm Halo}<2.0\times 10^{12}~{\rm M}_{\odot}$ and no massive ($M_{\rm Halo}>10^{12}~{\rm M}_{\odot}$) galaxy within $1$ Mpc at z=0z~\!\!=~\!\!0. These zoom-in systems are then resimulated with hydrodynamics and baryonic feedback turned on and increasing the resolution of all particles within $5R_{200}$.
@@
The key advantage of the DREAMS simulations is that they contain wide and systematic variations over the fiducial TNG astrophysical ($\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,\mathrm{high}}$) and cosmological parameters ($\Omega_{\rm M}$ and $\sigma_{8}$), as well as a range of initial conditions that sample rare environments and formation histories. The simulation parameters are varied according to a Sobol’ (1967) sequence where each simulation represents its own unique realization of TNG physics and cosmology (i.e., SB5 or “Sobol 5” in the parlance of the CAMELS simulations; Villaescusa-Navarro et al. 2021). Detailed motivation for the parameter ranges can be found in Rose et al. (2025a, their Section 2.1). In the following two subsections, we define each of the varied parameters as well as their ranges, as summarized in Table 1. Aside from the two cosmological parameters explicitly varied, the rest of the cosmology in DREAMS is fully consistent with Planck Collaboration et al. (2016): $H_{0}=100h~{\rm km\,s^{-1}\,Mpc^{-1}}$ where $h=0.6909$, and $\Omega_{b}=0.046$.
@@ 2.1.1 Parameter Variations
$\dot{E}_{\rm wind}\propto\eta v_{w}^{2}~.$ (5)
The impact of $\bar{e}_{w}$ is clear: increasing $\bar{e}_{w}$ raises the mass-loading factor, which increases the energy injection rate. Increases to $\kappa_{w}$, on the other hand, have a more subtle effect. By increasing $\kappa_{w}$, the wind speed increases, however, the mass-loading factor is proportionally decreased (since $\eta\propto v_{w}^{-2}$) keeping the energy injection rate fixed. Separate to the energy injection rate, the “efficacy” of feedback is also important. That is, whether launching less material faster has a strong er impact on the system than more material slower. Here we define “efficacy” in terms of the galaxy-scale consequences of feedback, such as reduced stellar mass growth and stronger baryon removal, instead of the instantaneous energy injection rate (which is fixed in the case of $\kappa_{w}$ variations). For the range of values probed in this work, increasing $\kappa_{w}$ increases the “efficacy”—by our definition—of feedback-driven winds (i.e., in this part of the TNG model parameter space it is more effective to drive slightly less material at slightly faster speeds). We show this effect explicitly in Section 3.3.
@@
Figure 1: Milky Way Mass Dark Matter Halos from the SB5 DREAMS CDM Suite. Projections of the dark matter density of a small fraction ($3\%$) of DREAMS halos. Each row contains variations of the five DREAMS parameters ($\bar{e}_{w}$, $\kappa_{w}$, $\epsilon_{f,\,\mathrm{high}}$, $\Omega_{\rm M}$, and $\sigma_{8}$, top-to-bottom, respectively; see Table 1 and Section 2.1.1). Each column represents the minimum, $20^{\rm th}$ percentile, $40^{\rm th}$ percentile, $60^{\rm th}$ percentile, $80^{\rm th}$ percentile, and maximum variation for each parameter, noting that in each row the other DREAMS parameters are also co-varied. The halos simulated here are all generated using different initial conditions (see Rose et al. 2025a). Thus, these simulations capture: (i) the impact of uncertainty in cosmology and astrophysics, (ii) the uncertainty in estimation of Milky Way halo mass, and (iii) intrinsic halo-to-halo variation.
2.2 Reconstruction of Dark Matter Density
The simulation dark matter is modeled as a collection of (large) discrete particles in TNG (see Section 2.1 for mass resolution). Since the goal of this paper is to study the dark matter profiles, we need to “reconstruct” a continuous density from these discrete particles. To obtain the density at each point in space, we distribute the mass of a single dark matter particle according to a smooth particle hydrodynamics kernel weighting scheme (e.g., Monaghan, 1992; Springel, 2010b). The density at a location in space, $s$, is computed as

### v2

@@ Abstract
In this work, we utilize a new suite of Milky Way-mass halos from the DREAMS Project, simulated with Cold Dark Matter (CDM), to quantify the influence of baryon feedback and intrinsic halo-to-halo variance on dark matter density profiles. Our suite of 1024 halos varies over supernova and black hole feedback parameters from the IllustrisTNG model, as well as variations in two cosmological parameters. We find that, for the DREAMS parameter variations, Milky Way-mass dark matter density profiles in the IllustrisTNG model are largely insensitive to astrophysics and cosmology variations, with the dominant source of scatter instead arising from halo-to-halo variance. However, most of the (comparatively minor) feedback-driven variations come from the changes to supernova prescriptions. By comparing to dark matter-only simulations, we find that the strongest supernova wind energies are so effective at preventing galaxy formation that the halos are nearly entirely collisionless dark matter. Finally, regardless of physics variation, all the DREAMS halos are roughly consistent with a halo contracting adiabatically from the presence of baryons, unlike models that have bursty stellar feedback. This work represents a step toward assessing the uncertainty in Milky Way dark matter profiles, with direct implications for dark matter searches where systematic uncertainty in the density profile remains a major challenge.
@@
I Introduction
Baryons make up only a small fraction of the total mass budget of the Universe (Planck Collaboration et al. 2016), yet they play a central role in driving galaxy evolution. Unlike the dark matter component, which is assumed to be cold and collisionless in the standard Cold Dark Matter (CDM) model, the baryonic component experiences hydrodynamical forces and energy dissipation. These interactions ultimately facilitate the condensation of gas at the centers of halos and the formation of the galaxies we observe (Fall & Efstathiou, 1980; Blumenthal et al., 1986). Once a galaxy is formed, baryons can impart feedback in their environments via various processes including supernovae, stellar winds, and active galactic nuclei (AGN) activity (e.g., Dekel & Silk, 1986; Larson, 1974; Somerville & Davé, 2015). Through interactions, dissipation, and feedback, baryons drive time-varying changes in the gravitational potential that shape galaxy evolution.
The role that baryons play can be clearly seen in the distribution of matter within a halo. Early dark matter-only (DMO) simulations showed that halos ubiquitously form a two-component power-law density profile known as the Navarro, Frenk, & White (1997; NFW) profile, such that
@@ I Introduction
Simulations which self-consistently model the baryonic component provide one possible resolution to these tensions (Sales et al., 2022). A cored inner density can be achieved using sufficiently “bursty” supernova feedback (Chan et al., 2015; Lazar et al., 2020; Mostow et al., 2024). Repetitively cycling between slow/smooth inflows followed by fast/impulsive outflows creates central potential fluctuations that disrupt the orbits of dark matter particles and lead to decreased central densities (Governato et al., 2010; Pontzen & Governato, 2012; Oñorbe et al., 2015; Jahn et al., 2023, although this effect is mass-dependent with it being most pronounced in dwarf galaxies). However, even at fixed physics prescriptions, the diversity of halo assembly histories introduces substantial halo-to-halo variance in the resulting dark matter profiles (e.g., Abadi et al., 2010; Duffy et al., 2010; Di Cintio et al., 2014; Chan et al., 2015; Fitts et al., 2017; Lazar et al., 2020; Farahi et al., 2022). Disentangling this intrinsic variance from systematic shifts due to baryonic physics is therefore critical for interpreting both simulations and observations. Moreover, the level to which feedback is generically bursty and its evolutionary consequences across different galaxy populations remains uncertain (see, e.g., work by Hartley & Ricotti, 2016; Bhagwat et al., 2024; Garcia et al., 2023, 2024a, 2024b, 2025a, 2025b, 2025c).
@@
Current cosmological simulations of galaxy evolution follow several prescriptions for modeling the physics of baryons (see, e.g., Vogelsberger et al., 2020; Crain & van de Voort, 2023; Feldmann & Bieri, 2025, for reviews). These prescriptions often rely on “subgrid” models for the numerical implementations of unresolved physics. For example, the interstellar medium (ISM) is often treated with an effective equation of state that sets the behavior of dense ($n_{\rm H}\gtrsim 0.1~{\rm cm}^{-3}$) star-forming gas (Springel & Hernquist, 2003; Schaye & Dalla Vecchia, 2008). Subgrid models require assumptions and oftentimes contain tunable parameters within the model. As a result, different simulation models make different predictions for the role of baryons and their associated feedback—even with relatively similar implementations (e.g., Chua et al., 2019, 2022; Garcia et al., 2024a, 2025a, 2025b). Some work has been done to quantify the variation in model predictions based on the input physics (e.g., Duffy et al., 2010; Torrey et al., 2014; Pillepich et al., 2018a; Chua et al., 2019, 2022; Font et al., 2020; Anbajagane et al., 2022; Kugel et al., 2023), but the number of model variations is usually few (on the order of $\sim 10$s of model variations). Recent simulation efforts such as CAMELS (Villaescusa-Navarro et al., 2021; Ni et al., 2023) and DREAMS (Rose et al., 2025a, hereafter R25a) have begun to more systematically explore the parameter space of well-tested simulation models via thousands of model variations. These efforts allow for quantification of parameter uncertainties while also providing large enough samples to begin quantifying the intrinsic halo-to-halo variance.
@@
One model that departs sharply from the standard approach is FIRE (Hopkins et al., 2014, 2018), which explicitly resolves the multiphase interstellar medium. While FIRE has its own subgrid assumptions and parameters, the treatment of the star-forming ISM is a marked improvement compared to the standard approach of an effective equation of state. One feature of the FIRE model is the production of strong, time-variable stellar feedback-driven winds (i.e., bursty feedback; Muratov et al., 2015), whereas models that adopt effective equation of state ISM treatment yield smoother, less variable stellar feedback. In the absence of this bursty feedback, the evolution of a dark matter halo is governed by the response of the dark matter to the baryons cooling to the center. This contraction can be modeled via the conservation of action integrals (e.g., Eggen et al., 1962; Blumenthal et al., 1986; Gnedin et al., 2004; Cautun et al., 2020): as the potential deepens due to baryon mass accumulation, the radius of the dark matter orbits decreases causing the total density in the center to increase. Feedback can partially offset or even reverse this contraction, but the degree depends sensitively on feedback (Brooks & Zolotov, 2014; Lovell et al., 2018; Hussein et al., 2025). Indeed, Hussein et al. (2025) show that FIRE-2 galaxies deviate much more from simple adiabatic contraction models than simulations with equation of state ISM prescriptions, though it remains unclear whether this owes specifically to burstiness or just the overall feedback strength.
@@
The rest of this paper is organized as follows. In Section II, we introduce the DREAMS simulations, their parameter variations, and halo selection. We then outline our method for reconstructing dark matter density fields, describe the neural network emulators used in this work, and discuss a scheme to account for how well our halos match observed galaxy scaling relations. In Section III, we discuss the dark matter density profiles and fit them with an analytic function. In Section IV, we compare the hydrodynamic simulations to DMO simulations, quantify the contraction the halos must have undergone due to the presence of baryons, and contrast our results to those from the bursty feedback FIRE model. In Section V, we state our conclusions.
II Methods
In this Section, we describe the DREAMS CDM suite (§II.1), describe how we reconstruct and fit the dark matter density profiles (§II.2), introduce our neural network emulation scheme (§II.3), and summarize the (Rose et al., 2025b, hereafter R25b) halo weighting scheme we apply to our sample (§II.4).
II.1 DREAMS Simulations
@@
Table 1: Astrophysical and Cosmological Parameter Variations. The astrophysical ($\bar{e}_{w}$, $\kappa_{w}$, $\epsilon_{f,\,\mathrm{high}}$) and cosmological ($\Omega_{\rm M}$, $\sigma_{8}$) parameters varied in the CDM DREAMS Milky Way-mass zoom-in simulations. We list the fiducial TNG value for each parameter, the variations used in these simulations (corresponding to a factor of 44 for $\bar{e}_{w}$ and $\epsilon_{f,\,\mathrm{high}}$, a factor of 22 for $\kappa_{w}$, and the two-standard-deviation uncertainty on Planck Collaboration et al. 2014 for $\Omega_{\rm M}$ and $\sigma_{8}$). We sample the unique set of parameters for each simulation according to a Sobol’ (1967) sequence. Finally, we note that $\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,\mathrm{high}}$ are sampled in logarithmic space, whereas $\Omega_{\rm M}$ and $\sigma_{8}$ are sampled in linear space. Detailed justifications for these parameter ranges can be found in R25a.
[…]
@@
The hydrodynamic simulations are based on the IllustrisTNG (hereafter simply TNG) physics model (Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018; Pillepich et al., 2018b; Springel et al., 2018; Pillepich et al., 2019; Nelson et al., 2019a, b). TNG is built on the moving Voronoi mesh code arepo (Springel, 2010a). The TNG model implements a wide range of astrophysical processes, including (but not limited to) feedback from stars, the growth of supermassive black holes, feedback from AGN, cosmological expansion of the universe, gravity, dark matter, and galactic winds; we refer the reader to Pillepich et al. (2018a) for a complete description of the TNG model.
@@
The DREAMS halos are selected using an iterative zoom-in process (described in detail in R25a, R25b). Briefly, the halos are selected initially from a low-resolution $(100~{\rm Mpc}/h)^{3}$ DMO simulation and are then resimulated with DMO zooms. The target halos of the intermediate zooms are selected to have masses of $5\times 10^{11}~{\rm M}_{\odot}<M_{\rm Halo}<2.0\times 10^{12}~{\rm M}_{\odot}$ and no massive ($M_{\rm Halo}>10^{12}~{\rm M}_{\odot}$) galaxy within $1$ Mpc at z=0z~\!\!=~\!\!0. These zoom-in systems are then resimulated with hydrodynamics and baryonic feedback turned on and increasing the resolution of all particles within $5R_{200}$.
@@
The key advantage of the DREAMS simulations is that they contain systematic variations over the fiducial TNG astrophysical ($\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,\mathrm{high}}$) and cosmological parameters ($\Omega_{\rm M}$ and $\sigma_{8}$), as well as a range of initial conditions that sample rare environments and formation histories. The simulation parameters are varied according to a Sobol’ (1967) sequence where each simulation represents its own unique realization of TNG physics and cosmology (i.e., SB5 or “Sobol 5” in the parlance of the CAMELS simulations; Villaescusa-Navarro et al. 2021). Detailed motivation for the parameter ranges can be found in R25a (their Section 2.1). In the following two subsections, we define each of the varied parameters as well as their ranges, as summarized in Table 1. Aside from the two cosmological parameters explicitly varied, the rest of the cosmology in DREAMS is fully consistent with Planck Collaboration et al. (2016): $H_{0}=100h~{\rm km\,s^{-1}\,Mpc^{-1}}$ where $h=0.6909$, and $\Omega_{b}=0.046$.
@@ II.1.1 Parameter Variations
$\dot{E}_{\rm wind}\propto\eta_{w}v_{w}^{2}~.$ (5)
The impact of $\bar{e}_{w}$ is clear: increasing $\bar{e}_{w}$ raises the mass-loading factor, which increases the energy injection rate. Increases to $\kappa_{w}$, on the other hand, have a more subtle effect. By increasing $\kappa_{w}$, the wind speed increases, however, the mass-loading factor is proportionally decreased (since $\eta_{w}\propto v_{w}^{-2}$) keeping the energy injection rate fixed. Separate from the energy injection rate, the “efficacy” of feedback is also important. That is, whether launching less material faster has a stronger impact on the system than launching more material slower. Here we define “efficacy” in terms of the galaxy-scale consequences of feedback, such as reduced stellar mass growth and stronger baryon removal, instead of the instantaneous energy injection rate (which is fixed in the case of $\kappa_{w}$ variations). For the range of values probed in this work, increasing $\kappa_{w}$ increases the “efficacy”—by our definition—of feedback-driven winds (i.e., in this part of the TNG model parameter space it is more effective to drive slightly less material at slightly faster speeds). We show this effect explicitly in Section III.3.
@@
Figure 1: Milky Way Mass Dark Matter Halos from the SB5 DREAMS CDM Suite. Projections of the dark matter density of a small fraction ($3\%$) of DREAMS halos. Each row contains variations of the five DREAMS parameters ($\bar{e}_{w}$, $\kappa_{w}$, $\epsilon_{f,\,\mathrm{high}}$, $\Omega_{\rm M}$, and $\sigma_{8}$, top-to-bottom, respectively; see Table 1 and Section II.1.1). Each column represents the minimum, $20^{\rm th}$ percentile, $40^{\rm th}$ percentile, $60^{\rm th}$ percentile, $80^{\rm th}$ percentile, and maximum variation for each parameter, noting that in each row the other DREAMS parameters are also co-varied.
II.2 Reconstruction of Dark Matter Density
The simulation dark matter is modeled as a collection of (large) discrete particles in TNG (see Section II.1 for mass resolution). Since the goal of this paper is to study the dark matter profiles, we need to “reconstruct” a continuous density from these discrete particles. To obtain the density at each point in space, we distribute the mass of a single dark matter particle according to a smooth particle hydrodynamics kernel weighting scheme (e.g., Monaghan, 1992; Springel, 2010b). The density at a location in space, $s$, is computed as